OpenStudy (anonymous):

The ABC Toy Company is creating two similar pieces for a board game. Find the value of y that makes the two pieces similar.

OpenStudy (anonymous):



OpenStudy (anonymous):

@wmj259

hartnn (hartnn):

corresponding sides of similar triangles are 'proportional' so if 4 became 8 what will 6 become ?

OpenStudy (anonymous):

Would it become 12?

OpenStudy (wmj259):

@LiquidLuck007 , Correct

hartnn (hartnn):

yes, 12 is correct :)

hartnn (hartnn):

mathematical to write that would be \(\Large \dfrac{GH}{JK}=\dfrac{KL}{HI}=\dfrac{GI}{JL}\)

OpenStudy (anonymous):

Oh wow thank you, easier than I thought it was.

OpenStudy (wmj259):

@LiquidLuck007 Look at my profile picture

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Haha, it fits :)

OpenStudy (anonymous):

Thank you both.

hartnn (hartnn):

wmj, good luck for your engineering :)

OpenStudy (anonymous):

:D